f(x)=10x^(2)+13x-3, g(x)=5x+1. find (f o g)(x)
Answer by ikleyn(52780)

.
f(x)=10x^(2)+13x-3, g(x)=5x+1. find (f o g)(x)
~~~~~~~~~~~~~~~~~~~~

To get (f o g)(x), we should replace x in the formula


    f(x) = 10x^(2)+13x-3


by 5x+1.  Doing it mechanically, we get


    (f o g)(x) = 10*(5x+1)^2 + 13*(5x+1) - 3 = 10*(25x^2 + 10x + 1) + 65x + 13 - 3 = 

               = 250x^2 + 100x + 10 + 65x + 10 = 250x^2 + 165x + 20.    ANSWER

Solved, answered and explained.